OpenStudy (anonymous):

will give medal :DD 1.Factor completely a2 - a - 20 2.factor completely x2 + 13x + 36 3.Factor completely b2 + 12b + 32 4.factor completely x2 - 5x - 14

OpenStudy (anonymous):

20 factors 4,5 a^2-5a+4a-20=0=>a(a-5)+4(a-5)=0=>(a+4)(a-5)=0 similarly solve for others 36 factors are (6,6),(9,4),(12,3) 32 factors are (8,4) 14 factors are (2,7)

OpenStudy (anonymous):

:$ the answer? lol

OpenStudy (solomonzelman):

if you need just answers you can use online calculator. If you want to learn, then this is the right place....